OpenStudy (anonymous):

The graph of f(x) = 2^x + 1 is shown below. Explain how to find the average rate of change between x = 1 and x = 3.

OpenStudy (anonymous):

Can you tell me what f(x) is when x=1 and when x=3?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

it is simple, we have to apply this formula: \[r = \frac{{f\left( 3 \right) - f\left( 1 \right)}}{{3 - 1}}\]

OpenStudy (michele_laino):

now, we have: \(f(3)= 2^3+1=...\)

OpenStudy (anonymous):

ok:)

OpenStudy (michele_laino):

please continue, I will guide you to final answer

OpenStudy (michele_laino):

\(f(3)=...?\)

OpenStudy (anonymous):

1?

OpenStudy (michele_laino):

hint: \[f\left( 3 \right) = {2^3} + 1 = 8 + 1 = ...?\] since \(2^3=8\)

OpenStudy (anonymous):

https://www.desmos.com/calculator/d7rsqv1ntm \ This is what the graph looks like and isn't it asking us to find the slope then??

OpenStudy (michele_laino):

yes! Better is that we have to find an average rate, because, for this function, the term "slope" alone has no meaning

OpenStudy (anonymous):

ok:) So what do we do?

OpenStudy (michele_laino):

please complete: \(f(3)=8+1=...?\)

OpenStudy (anonymous):

9

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

now, we have to compute this: \(f(1)=2^1+1=2+1=...\)

OpenStudy (anonymous):

3

OpenStudy (michele_laino):

that's right! so after a substitution, into my formula above, we get: \[r = \frac{{f\left( 3 \right) - f\left( 1 \right)}}{{3 - 1}} = \frac{{9 - 3}}{{3 - 1}} = ...?\]

OpenStudy (anonymous):

6/2=3

OpenStudy (michele_laino):

that's right!! :)

OpenStudy (anonymous):

now what?

OpenStudy (michele_laino):

we have completed the requested solution

OpenStudy (anonymous):

did we? Ok so the average rate of change between x=1 and x=3 is 3?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thanks!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

Can I ask one more??

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

A biologist is comparing the growth of a population of flies per week to the number of flies a lizard will consume per week. She has devised an equation to solve for which day (x) the lizard would be able to eat the entire population. The equation is 3x = 5x − 1. Explain to the biologist how she can solve this on a graph using a system of equations

OpenStudy (anonymous):

I got 2,9 as the answer:)

OpenStudy (michele_laino):

I think that the requested system is like below: \[\left\{ \begin{gathered} y = 3x \hfill \\ y = 5x - 1 \hfill \\ \end{gathered} \right.\] since we need to introduce a new variable \(y\)

OpenStudy (anonymous):

yes:) but it's actually 3^x

OpenStudy (anonymous):

sorry :)

OpenStudy (anonymous):

y=3^x<---log function y=5x-1

OpenStudy (michele_laino):

ok! then we have to solve this system: \[\left\{ \begin{gathered} y = {3^x} \hfill \\ y = 5x - 1 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

\[\huge \left\{ \begin{gathered} y = {3^x} \hfill \\ y = 5x - 1 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

https://www.desmos.com/calculator/mtmjle2eyr this is the graph

OpenStudy (michele_laino):

so, we have to draw both graphs of the functions \(\Large y=3^x\) and \(\Large y=5x-1\)

OpenStudy (anonymous):

https://www.desmos.com/calculator/mtmjle2eyr Yes:)

OpenStudy (michele_laino):

correct! Your answer is right!

OpenStudy (anonymous):

2,9? Yay!

OpenStudy (michele_laino):

yes! Since we are searching for a number of days \(x\), so \(x\) has to be a whole number

OpenStudy (anonymous):

cool! I have this LAST question that I'm really confused on :(

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

John has taken out a loan for college. He started paying off the loan with a first payment of $100. Each month he pays, he wants to pay back 1.1 times as the amount he paid the month before. Explain to John how to represent his first 20 payments in sigma notation. Then explain how to find the sum of his first 20 payments, using complete sentences. Explain why this series is convergent or divergent.

OpenStudy (michele_laino):

those payments starts with $100, and at the second day the payment is \(1.1 \cdot 100\) at the third day it is \(1.1 \cdot (1.1 \cdot 100))=1.1^2 \cdot 100\), at the fourth day such payment is \(1.1 (1.1^2 \cdot 100)=1.1^3 \cdot 100\) am I right?

OpenStudy (anonymous):

yes:)

OpenStudy (michele_laino):

sorry! I meant months not days

OpenStudy (anonymous):

It's okay:)

OpenStudy (anonymous):

now what?

OpenStudy (michele_laino):

so at the fourth month the total payment is: \[\Large \begin{gathered} 100 + 1.1 \cdot 100 + {1.1^2} \cdot 100 + {1.1^3} \cdot 100 = \hfill \\ \hfill \\ = \sum\limits_{x = 1}^4 {{{1.1}^{x - 1}} \cdot 100} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

arent we suppose to find his first 20 payments?

OpenStudy (anonymous):

so then the 4 becomes 20 at the top?

OpenStudy (michele_laino):

If we consider the initial $100 as the first payment then yes, we have to replace 4 with 20

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

so what would be the answer?

OpenStudy (michele_laino):

so, in that case total payment is: \[\huge P = \sum\limits_{x = 1}^{20} {{{1.1}^{x - 1}} \cdot 100} \]

OpenStudy (michele_laino):

now, we have to see if such series is convergent

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

I think it's convergent

OpenStudy (michele_laino):

if the upper limit is \(+ \infty\) and not \(20\), then such series is divergent

OpenStudy (anonymous):

yeah, but the upper limit is 20. And now we have tto solve it

OpenStudy (anonymous):

5727.23?

OpenStudy (michele_laino):

yes! In order to do that, we can rewrite such series as below: \[\Large P = \sum\limits_{x = 1}^{20} {{{1.1}^{x - 1}} \cdot 100} = \frac{{100}}{{1.1}} \cdot \sum\limits_{x = 1}^{20} {{{1.1}^x}} \]

OpenStudy (michele_laino):

please wait, I'm checking your result...

OpenStudy (anonymous):

ok:)

OpenStudy (michele_laino):

I got this: \[\Large P = \sum\limits_{x = 1}^{20} {{{1.1}^{x - 1}} \cdot 100} = \frac{{100}}{{1.1}} \cdot \sum\limits_{x = 1}^{20} {{{1.1}^x}} \cong 5,727.5\]

OpenStudy (anonymous):

yay! Thank-you so much<3

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

i need help on 1 more:)

OpenStudy (anonymous):

A field test for a new exam was given to randomly selected seniors. The exams were graded, and the sample mean and sample standard deviation were calculated. Based on the results, the exam creator claims that on the same exam, nine times out of ten, seniors will have an average score within 4% of 70%. Is the confidence interval at 90%, 95%, or 99%? What is the margin of error? Calculate the confidence interval and explain what it means in terms of the situation.

OpenStudy (anonymous):

i think it's 90%?

OpenStudy (michele_laino):

I'm very sorry, I don't know this answer :(

OpenStudy (michele_laino):

why do you think it is 90%

OpenStudy (anonymous):

um well I guessed :) Can you google the answer?

OpenStudy (michele_laino):

I don't know. I think nine times out of ten, it is equivalent to 9/10, nevertheless I'm not sure, please ask to another helper

OpenStudy (michele_laino):

@dan815 please can you help here?

OpenStudy (anonymous):

Thanks for trying:)

OpenStudy (michele_laino):

:)

OpenStudy (dan815):

yeah 90% confidence interval

OpenStudy (anonymous):

bUT what would the answer be??